# Towards Controllable Speech Synthesis in the Era of Large Language Models: A Survey

### v1

@@ VI-A Challenges
Controllable granularity. A critical challenge in controllable TTS is determining what aspects of speech should be controlled and how to control speech characteristics at a specific granularity. Different applications require varying levels of control granularity. For instance, audiobook narration may need sentence-level control of emotion, while conversational AI like ChatGPT may require word or phoneme-level control over prosody. Moreover, the emotion, prosody, and other characteristics of human speech are often intricately intertwined and can manifest across varying levels of granularity. Additionally, achieving fine-grained control requires high-resolution annotations and sophisticated models capable of handling subtle variations without compromising synthesis quality.
@@
Feature extraction and representation. Achieving fully controllable TTS needs good feature disentanglement. Accurately extracting meaningful and disentangled speech features like pitch contours, energy patterns, emotion variation, and prosodic elements from training data is difficult. The reason is that speech features are interdependent and context-sensitive, making it hard to isolate specific attributes for control. For example, altering pitch often affects prosody, emotion, and naturalness to some extent. To tackle this, several methods [269, 270, 271] utilize pre-trained models for different speech recognition tasks (e.g., pitch, energy, and duration prediction, gender classification, age estimation, and speaker verification) to supervise feature extraction. For example, NaturalSpeech3 [14] factorizes speech into separate feature subspaces to capture different speech attributes.
@@
Scarcity of datasets. High-quality, diverse, and appropriately annotated datasets are essential for training controllable TTS systems. However, such datasets are scarce and difficult to construct. To achieve controllable TTS, training data must encompass a wide range of styles, emotions, accents, and prosodic variations to enable versatile control because limited diversity in datasets can restrict the model’s ability to generalize across unseen styles or emotions. Although there are some TTS datasets, such as LibriTTS [272], Gigaspeech [258], and TextrolSpeech [247], their diversity is still not enough for fully controllable TTS due to the lack of corpus of diverse content such as comedies, thrillers, cartoons, etc. Constructing large-scale datasets with rich diversity is also expensive and time-consuming.
@@
Another obstacle is that creating datasets with fine-grained, attribute-specific annotations is labor-intensive and costly. Besides, manual annotation of speech attributes requires expert knowledge and is prone to inconsistencies and errors, particularly for subjective qualities like emotion. Currently, most datasets provide only coarse labels, such as gender, age, or a limited range of emotions. While some datasets, such as SpeechCraft [262], include natural language descriptions of speech attributes, no existing dataset offers fine-grained variations and annotations within the speech of the same speakers. Available datasets for controllable TTS are summarized in Table V.
@@
Generalization ability. The ability of a TTS system to generalize effectively is crucial for producing natural, high-quality speech across a wide range of conditions, such as unseen speakers, languages, or topics. However, achieving robust generalization remains a significant challenge for modern TTS methods due to various factors. Zero-shot controllable TTS [78, 92] aims to synthesize speech for unseen speakers with various speech customization such as emotion using minimal reference audio, which can offer flexibility for personalized voice generation. However, it faces significant challenges, including capturing unique speaker characteristics from limited data, accurately reproducing prosody and style, and disentangling speaker identity from other audio attributes like emotion or noise.
@@
Multilingual generalization [250, 273] in TTS refers to the ability to synthesize natural and intelligible speech across multiple languages, including those not seen during training. This capability is essential for applications like cross-lingual communication, multilingual virtual assistants, and speech synthesis for low-resource languages [274]. Multilingual generalization still faces many challenges such as linguistic diversity and mismatch and the scarcity of data. Cross-lingual speaker generalization is another hurdle, as preserving speaker identity across languages can lead to artifacts.
@@
Domain adaptation [275] in TTS refers to tailoring a pre-trained TTS model to generate speech for a specific domain or context, such as medical terminology and conversational speech. One challenge is that many specialized domains lack sufficient high-quality annotated data for fine-tuning. Besides, adapting prosody, intonation, and speaking style to match domain-specific requirements such as comic dialogue is complex. Failing to capture domain-specific nuances can make speech sound unnatural or inconsistent with the target context.
@@
High latency is a major issue, as existing controllable TTS models [78, 101, 102, 103] often necessitate autoregressive processes to synthesize speech. The inference time of these models ranges from several to tens of seconds. This can be especially problematic for real-time applications like live broadcasting or interactive systems. Additionally, the challenge of balancing granularity and efficiency arises, as finer controls demand higher-resolution data and more precise models, leading to increased resource requirements and inefficient training and inference.
Another major obstacle lies in the trade-off between model complexity and performance. State-of-the-art controllable TTS systems often rely on large neural networks such as LLMs with billions of parameters, which provide superior naturalness and expressiveness but demand significant computational resources. Simplifying these architectures can lead to quality degradation, including artifacts, unnatural prosody, or limited expressiveness. Therefore, designing light-weight controllable TTS models is significantly tricky.
@@
Figure 4: A summary of current major challenges for fully controllable TTS.
@@ VI-B Future Directions
In this survey, we conduct a comprehensive investigation and analysis of existing TTS methods, particularly on controllable TTS technologies.  While these methods show great potential in real-world applications, there are still some limitations that need to be addressed.  Based on our observations, we outline several promising future directions as follows:
@@
Fine-grained speech synthesis by natural language description. Using natural language description to synthesize human speech with fine-grained control over various audio attributes is currently underexplored. Most of the existing works can only control a fixed number of attributes of the synthesized speech. Although a few works show great control of emotion, timbres, pitch, gender, and styles, e.g., VoxInstruct [103] and CosyVoice [17], they can frequently synthesize unwanted speech clips. Users need to synthesize multiple times to get satisfactory speech.
@@
Fine-grained speech editing by natural language description. Speech or audio editing has been studied for a long time. However, existing methods usually train conditional models and adjust a fixed number of conditional inputs to modify the attributes of synthesized speech, thus lacking fine-grained manipulations [95, 94]. Therefore, how to learn disentangled speech representations for speech attributes while supporting editing by using natural language description is worthy of investigation.
@@
Expressive multi-modal speech synthesis. Synthesizing speech from multi-modal data such as texts, images, and videos is an appealing research topic due to its various applications in the industry such as storytelling, filming, and gaming. Although there are several related works on this task [6, 24, 276, 277], few of them can fully extract useful information from multi-modal data. Particularly, synthesizing engaging speech and expressive voiceover for complex visual content sees great opportunities in the future.
@@
Natural and emotional conversational TTS. Speech conversational TTS have come out for several decades but remained as cascaded systems for a long time and cannot generate natural and emotional speech. These systems are not context-aware, making the synthesized speech sound robotic. With the advent of LLMs, existing TTS technologies were directly introduced by simply synthesizing speech from the text generated by LLMs [33]. However, context-aware conversational TTS with rich emotion and good naturalness has not been well studied.
@@
Zero-shot long speech synthesis with emotion consistency. Zero-shot TTS emerged in recent years to achieve voice cloning and speech style imitation without fine-tuning, making them more practical in real scenarios [78, 194, 17]. However, synthesizing long speech with rich emotion and style variation in a zero-shot setting remains challenging due to the lack of rich speech information in short reference audio clips. Addressing this issue will make a big step towards fully controllable zero-shot TTS.
@@
Efficient TTS by natural language description. Synthesizing speech with natural language description usually involves training large language encoders and bridge nets between the two modalities which can bring about much more computation overhead compared to previous TTS methods. The inference time is also relatively slow, e.g., existing methods usually take tens of seconds to synthesize a short speech audio clip of less than 10 seconds [17, 104]. Therefore, efficient text and speech modeling and interaction is critical for natural language description-based TTS systems.
VII Discussion
VII-A Impacts of Controllable TTS
Applications. Controllable TTS systems allow fine-grained manipulation of speech attributes such as pitch, emotion, speaking rate, and style, enabling a wide range of applications across industries. One major application is in virtual assistants and customer support systems, where controllable TTS ensures tailored and context-aware responses. For instance, a virtual assistant can speak in a calm tone for technical support but switch to an enthusiastic tone when presenting promotional offers, enhancing user experience.
In the entertainment industry, controllable TTS is invaluable for creating dynamic voiceovers, audiobooks, and gaming characters. It enables precise adjustments in tone and delivery, allowing audiobooks to reflect character emotions and gaming characters to exhibit personality traits that align with their roles. Similarly, in education, TTS systems can adapt speaking styles to suit different learners, such as adopting a slow, clear tone for language learning or an engaging, storytelling style for children’s content.
Controllable TTS is also transformative in assistive technologies, where it can generate speech that reflects the user’s intended emotion or personality. This is particularly impactful for individuals with speech impairments, enabling more expressive and natural communication.
@@ VII-A Impacts of Controllable TTS
In content localization, controllable TTS systems adjust speech characteristics to suit regional and cultural preferences, ensuring a natural fit for global audiences. Additionally, it finds applications in human-computer interaction research, enabling adaptive dialogue systems that modify speech dynamically based on user mood or environment. By offering this flexibility, controllable TTS systems enhance accessibility, personalization, and engagement across various domains.
@@
Deepfakes. A deepfake is a type of synthetic media in which a person in an existing image, video, or audio recording is replaced with someone else’s likeness or voice. This technology uses deep learning, particularly GANs [278], to create highly realistic, but fabricated, content. While deepfakes are most commonly associated with video manipulation, such as face swapping [279], they can also apply to audio, enabling the creation of synthetic speech that mimics a specific person’s voice, which is well known as voice cloning.
Voice cloning, especially few-shot [280] and zero-shot TTS [85, 78], poses a significant threat to systems that rely on voice authentication, such as banking, customer service, and other identity verification processes. With a convincing clone of someone’s voice, attackers could potentially impersonate individuals to gain unauthorized access to sensitive information or accounts.
@@
VII-B Limitation of this Survey
Although we conduct a comprehensive survey of controllable TTS in this paper, there are some limitations we want to address in the future: 1) A unified benchmark method will be provided to evaluate controllable TTS methods. 2) Detailed analysis and control strategies of each specific speech attribute will be provided in an updated version of this paper. 3) The methodologies for feature disentanglement are crucial for controllable TTS but are not adequately discussed.
@@ VIII Conclusion
In this survey paper, we first elaborate on the general pipeline for controllable TTS, followed by a glimpse of uncontrollable TTS methods. Then we comprehensively review existing controllable methods from the perspectives of model architectures and control strategies. Popular datasets and commonly used evaluation metrics for controllable TTS are also summarized in this paper. Besides, the current challenges are deeply analyzed and the promising future directions are also pointed out. To the best of our knowledge, this is the first comprehensive survey for controllable TTS.
Writing a comprehensive survey is a challenging task and an iterative process. Hence, we will regularly update this survey to offer researchers and practitioners a continuously evolving resource for understanding controllable speech synthesis.

### v2

@@ VI-A Challenges
Controllability. A critical challenge in controllable TTS is determining what aspects of speech should be controlled and how to control speech characteristics at a specific granularity. Different applications require varying levels of control granularity. For instance, audiobook narration may need sentence-level control of emotion, while conversational AI like ChatGPT may require word or phoneme-level control over prosody. Moreover, the emotion, prosody, and other characteristics of human speech are often intricately intertwined and can manifest across varying levels of granularity. Additionally, achieving fine-grained control requires high-resolution annotations and sophisticated models capable of handling subtle variations without compromising synthesis quality.
@@
Feature Disentanglement and Representation. Achieving fully controllable TTS needs good feature disentanglement. Accurately extracting meaningful and disentangled speech features like pitch contours, energy patterns, emotion variation, and prosodic elements from training data is difficult. The reason is that speech features are interdependent and context-sensitive, making it hard to isolate specific attributes for control. For example, altering pitch often affects prosody, emotion, and naturalness to some extent. To tackle this, several methods [303, 304, 305] utilize pre-trained models for different speech recognition tasks (e.g., pitch, energy, and duration prediction, gender classification, age estimation, and speaker verification) to supervise feature extraction. For example, NaturalSpeech3 [14] factorizes speech into separate feature subspaces to capture different speech attributes.
@@
Scarcity of Datasets. High-quality, diverse, and appropriately annotated datasets are essential for training controllable TTS systems. However, such datasets are scarce and costly to construct. In addition, training data must encompass a wide range of styles, emotions, accents, and prosodic variations to enable versatile control because limited diversity in datasets can restrict the model’s ability to generalize across unseen styles or emotions. Although there are some large-scale datasets, such as LibriTTS [306], Gigaspeech [285], and TextrolSpeech [243], their diversity is still not enough for fully controllable TTS due to the lack of corpora for diverse content and scenarios such as comedies, thrillers, and cartoons. Constructing these large-scale datasets with rich diversity is expensive and time-consuming.
@@
Another obstacle is that creating datasets with fine-grained, attribute-specific annotations is labor-intensive and costly. In addition, manual annotation of speech attributes requires expert knowledge and is prone to inconsistencies and errors, particularly for subjective attributes like emotion. Currently, most datasets provide only coarse labels, such as gender, age, or a limited range of emotions. While some datasets, such as SpeechCraft [291] and Parler-TTS [273], include natural language descriptions of speech attributes, no existing dataset offers fine-grained variations and annotations within the speech of the same speakers. Publicly available datasets for controllable TTS are summarized in Table VI.
@@
Generalization Ability. The ability of a TTS system to generalize effectively is crucial for producing natural, high-quality speech across a wide range of conditions, such as unseen speakers, languages, and topics. However, achieving robust generalization remains a significant challenge for modern TTS methods due to various factors. Zero-shot controllable TTS [78, 92] aims to synthesize speech for unseen speakers with various speech customizations, such as emotion, using minimal reference audio, which can offer flexibility for personalized voice generation. However, it faces significant challenges, including capturing unique speaker characteristics from limited data, accurately reproducing prosody and style, and disentangling speaker identity from other attributes such as emotion and background noise.
@@
Multilingual & low-resource generalization [253, 307] in TTS refers to the ability to synthesize natural and intelligible speech across multiple languages, including those not seen during training. This capability is essential for applications like cross-lingual communication, multilingual virtual assistants, and speech synthesis for low-resource languages [308]. Multilingual generalization still faces many challenges, such as linguistic diversity, mismatch, and the scarcity of data. Cross-lingual speaker generalization is another hurdle, as preserving speaker identity across languages can lead to artifacts.
@@
Domain adaptation [309] in TTS refers to tailoring a pre-trained TTS model to generate speech for a specific domain or context, such as medical terminology and debate. One challenge is that many specialized domains lack sufficient high-quality annotated data for fine-tuning. In addition, adapting prosody, intonation, and speaking style to match domain-specific requirements such as comic dialogue is complex. Failing to capture domain-specific nuances can make speech sound unnatural or inconsistent with the target context.
@@
High latency is a major issue, as existing controllable TTS models [78, 101, 102, 103] often necessitate autoregressive processes to synthesize speech. The inference time for these models can range from several seconds to tens of seconds for a short audio clip (e.g., 5 to 10 seconds), and they do not support streaming synthesis. This can be particularly problematic for real-time applications like live broadcasting or interactive systems. Additionally, balancing granularity and efficiency presents a challenge. Finer controls require higher-resolution data and more precise modeling, which increases resource demands and results in inefficient training and inference.
Another major obstacle lies in the trade-off between model complexity and performance. SOTA controllable TTS systems often rely on LLMs with billions of parameters, which provide superior naturalness and expressiveness but demand significant computational resources. Simplifying these architectures can lead to quality degradation, including artifacts, unnatural prosody, or limited expressiveness. Therefore, designing light-weight controllable TTS models is significantly tricky.
@@
§$\S$§VI-A
Figure 5: A summary of current challenges for fully controllable TTS.
@@ VI-B Future Directions
In this survey, we conduct a comprehensive investigation and analysis of existing TTS methods, particularly on controllable TTS technologies. While these methods show great potential in applications, numerous limitations and challenges still need to be addressed. Hence, we are still in the early stage of controllable speech synthesis. Based on our observations, we outline several promising future directions as follows:
@@
Instruction-Guided Fine-grained Speech Synthesis. Using natural language description to synthesize human speech with fine-grained control over various speech attributes is currently underexplored. Most of the existing works can only control a fixed number of attributes of the synthesized speech. Although a few works show great control of emotion, timbres, pitch, gender, and styles, e.g., VoxInstruct [103] and CosyVoice [17], they can frequently synthesize unwanted speech clips inconsistent with user instructions. Users often need to synthesize multiple times to get satisfactory speech.
@@
Instruction-Guided Fine-grained Speech Editing. Speech or audio editing has been studied for a long time. However, existing methods usually train conditional models and adjust a fixed number of conditional inputs to modify the attributes of synthesized speech, thus lacking fine-grained manipulations [95, 94]. Therefore, how to learn disentangled speech representations for speech attributes while supporting instruction-guided editing is worthy of investigation.
@@
Expressive Multi-modal Speech Synthesis. Synthesizing speech from multi-modal data such as texts, images, and videos is an appealing research topic due to its various applications in the industry, such as storytelling, filming, and gaming. Although there are several related works on this task [6, 24, 310, 311], they are still limited in extracting the required information from multimodal data. Particularly, synthesizing engaging speech and expressive voiceover for complex visual content sees great opportunities in the future.
@@
Natural and Emotional Conversational TTS. Conversational TTS has been studied for several decades, but remained as cascaded systems for a long time, limiting its ability to generate natural and expressive speech. These systems are not context-aware, making the synthesized speech sound robotic. With the advent of LLMs, existing TTS technologies are directly introduced to synthesize speech using discrete speech tokens [33, 34]. However, context-aware conversational TTS with rich emotion and naturalness has not been well studied.
@@
Zero-shot Long Speech Synthesis with Emotion Consistency. Zero-shot TTS is capable of voice cloning and speech style imitation without fine-tuning, making it practical in real scenarios [78, 229, 17]. However, synthesizing long speech with rich emotion and style variation in a zero-shot setting remains challenging due to the limited information in short reference audio clips. Addressing this issue will make a big step towards personalized zero-shot TTS.
@@
Instruction-Guided Efficient TTS. Synthesizing speech with user instructions usually involves training language model-based codecs and bridge nets between different modalities, leading to much more computation overhead than previous TTS methods. The inference time is also relatively long, e.g., existing instruction-guided methods usually take tens of seconds to synthesize a short speech of less than 10 seconds [17, 104]. Therefore, efficient text and speech modeling is critical for instruction-guided TTS systems.
VII Impacts of Controllable TTS
VII-A Applications
Controllable TTS enables fine-grained manipulation of speech attributes like pitch, emotion, and style, making it valuable across industries. For example, in virtual assistants and customer support, controllable TTS ensures context-aware responses, such as a calm tone for technical help or an enthusiastic pitch for promotions. In entertainment, it enhances voiceovers, audiobooks, and gaming characters by adjusting tone and delivery to match emotions and personalities. Education benefits from adaptive TTS, like slow, clear articulation for language learning or engaging narration for children.
@@ VII-A Applications
For assistive technologies, controllable TTS empowers individuals with speech impairments to express emotions naturally. In content localization, it adapts speech to cultural preferences, ensuring a seamless experience for global audiences. Additionally, in human-computer interaction, it enables adaptive dialogue systems that adjust speech based on user mood or environment. By offering flexibility and expressiveness, controllable TTS enhances accessibility, personalization, and engagement across diverse applications.
VII-B Security Issues
@@
One major security issue brought by controllable TTS is deepfakes. A deepfake is a type of synthetic media in which a person in an existing image, video, or audio recording is replaced with someone else’s likeness or voice. This technology uses deep learning, particularly GANs [312], to create highly realistic but fabricated content. While deepfakes are most commonly associated with video manipulation, such as face swapping [313], they can also be applied to audio, enabling the creation of synthetic speech that mimics a specific person’s voice, which is well known as voice cloning. Voice cloning, especially few-shot [314] and zero-shot TTS [85, 78], poses a significant threat to systems that rely on voice authentication, such as banking, customer service, and other identity verification processes, allowing attackers to impersonate individuals to gain unauthorized access to sensitive information or accounts.
Another issue is adversarial attacks [315]. Attackers can manipulate pitch, prosody, and phoneme duration to fool automatic speech recognition and speaker verification. It also poses risks in data poisoning [316], where adversarially modified samples degrade model performance. On the other hand, controllable TTS can strengthen adversarial defenses by generating diverse speech variations for adversarial training and counter-adversarial augmentation [317].
@@ VIII Conclusion
In this survey, we have first elaborated on the general pipeline for controllable TTS, followed by a glimpse of uncontrollable TTS methods from the perspective of local and global speech modeling. Then, we have comprehensively reviewed existing controllable methods from the perspectives of model architectures and control strategies. Popular datasets and commonly used evaluation metrics for controllable TTS were also summarized in this paper. Besides, the current challenges were deeply analyzed, and the promising future directions were also pointed out. To the best of our knowledge, this is the first comprehensive survey for controllable TTS.